### v1

# Transfer Learning for Neutrino Scattering: Domain Adaptation with GANs

Abstract
We utilize transfer learning to extrapolate the physics knowledge encoded in a Generative Adversarial Network (GAN) model trained on synthetic charged-current (CC) neutrino–carbon inclusive scattering data. This base model is adapted to generate CC inclusive scattering events (lepton kinematics only) for neutrino–argon and antineutrino–carbon interactions. Furthermore, we assess the effectiveness of transfer learning in re-optimizing a custom model when new data comes from a different neutrino–nucleus interaction model. Our results demonstrate that transfer learning significantly outperforms training generative models from scratch. To study this, we consider two training data sets: one with 10,000 and another with 100,000 events. The models obtained via transfer learning perform well even with smaller training data. The proposed method provides a promising approach for constructing neutrino scattering event generators in scenarios where experimental data is sparse.
I Introduction
Significant experimental efforts have been devoted to studying (anti)neutrino–nucleus interactions [1, 2] in the energy range relevant for next-generation neutrino oscillation experiments, such as Hyper-Kamiokande [3] and DUNE [4]. In parallel, theoretical models describing these interactions have been developed [2]. The outcomes of both experimental and theoretical advances are incorporated into Monte Carlo (MC) event generators, which simulate (anti)neutrino–nucleus collisions under realistic conditions [5, 6, 7, 8, 9, 10].
MC generators are often tuned to reproduce experimental observations, relying on adjustable parameters that are fitted using available data [11]. However, this tuning process cannot fully compensate for the fundamental limitations of the underlying models, especially those relying on complex approximations, such as nuclear modeling. Consequently, there is a growing interest in alternative approaches to traditional MC event generation—methods that can learn directly from experimental data and dynamically refine their predictions. Among the most promising directions is the development of artificial intelligence (AI)-driven tools for simulating particle collisions [12, 13, 14].
In this paper, we present results from our development of generative deep learning models designed to simulate neutrino and antineutrino interactions with nuclei. Specifically, we focus on generative adversarial networks (GANs) [15], which have recently gained much interest in high-energy and nuclear physics [16, 17, 18, 19, 20, 21, 22].
Our previous paper [23] presented one of the first applications of GAN models for simulating neutrino collisions with nuclei. Additionally, Refs. [24, 25] provide further examples of using generative deep learning techniques in the simulation of neutrino-nucleus interactions, although in these studies, the method of normalizing flows is employed.
In our previous work, we investigated the properties of GAN models trained to reproduce synthetic inclusive and quasi-elastic charged-current (CC) muon neutrino scattering on carbon nucleus data obtained using the NuWro MC generator. However, our model only generated the kinematics of the charged muon for a given input neutrino energy. In this paper, we explore a technique known as domain adaptation, a form of transfer learning (TL) in which a model trained on one domain (the source domain) is adjusted to perform well on a different but related domain (the target domain). This technique is particularly promising when data in the target domain is scarce.
We previously explored a similar concept in Ref. [26], where a model trained on inclusive electron-carbon scattering data was adapted to predict electron scattering on helium-3, lithium, oxygen, aluminum, calcium, and iron. In this paper, we begin with a baseline model trained on inclusive CC muon neutrino scattering off the carbon nucleus. We then adapt this model to generate predictions for three different processes: muon neutrino scattering off the argon nucleus (relevant for the DUNE experiment), muon-antineutrino scattering off the carbon nucleus, and muon neutrino scattering off the carbon nucleus using data obtained from a different version and configuration of the NuWro generator than that used for training the baseline model.
To evaluate the effectiveness of the TL technique, we compare the results from our domain adaptation models with predictions from models that were trained from scratch. Our findings demonstrate that models employing domain adaptation yield more accurate predictions than those trained from scratch, highlighting the potential of this technique in situations with limited information.
This paper is structured as follows. In Sec. II, key features of the network and its training are explained. In Sec. IV we present the results of the training on multiple neutrino interaction models. Finally, in Sec. V we summarize the paper. In Appendix A, we discuss the properties of the loss function we used in our analysis.
II GAN framework
II.1 Kinematics
Our goal is to obtain a model that generates the kinematic characteristics of the charged muon produced in the inclusive CC $\nu_{\mu}$-nucleus and $\overline{\nu}_{\mu}$-nucleus scattering. We consider the same generator and discriminator models as in our previous paper [23].
The generator for given normalized (anti)neutrino energy $E_{\nu}^{\prime}$ gives as an output two kinematic variables characterizing the muon, namely, $E_{\mu}^{\prime}$ and $\theta^{\prime}$, where
$E_{\nu}^{\prime}=2\frac{E_{\nu}-E_{\nu,min}}{E_{\nu,max}-E_{\nu,min}}-1,$ (1)
$E_{\nu}$ is the (anti)neutrino energy, $E_{\nu,min}=0.3$ GeV, $E_{\nu,max}=10$ GeV,
$E_{\mu}^{\prime}$ $=$ $2\sqrt{1-\frac{E_{\mu}-m_{\mu}}{\Delta E}}-1,\quad\Delta E=E^{-}_{\nu}-m_{\mu}$ (2)
$\theta^{\prime}$ $=$ $2\sqrt{\frac{\theta}{\pi}}-1,$ (3)
and $m_{\mu}$ is the muon mass, $\theta$ is the muon scattering angle relative to the neutrino beam direction.
The choice of the proxy prime variables is discussed in the Ref. [23]. The values $E_{\nu,min}$ and $E_{\nu,max}$ correspond to minimal and maximal neutrino energy generated by the model, respectively. In this paper, we consider a slightly different value of $E^{-}_{\nu}$ than in the previous study [23]. While the previous boundary prevented the production of non-physical events, they turned out to be too strong for applications discussed in the present studies, namely, $\nu_{\mu}$-argon scattering, or changes in physical parameters in the simulation. Since in experimental data we expect not to follow the same physics as the NuWro simulation, the constraint has been relaxed to $E^{-}_{\nu}=E_{\nu}-4$ MeV.
In our previous studies, we optimized the model using $\nu_{\mu}$-carbon ($^{12}C$) scattering data produced with NuWro ver. 21.09 generator. In the present analysis, we generate data using the same version of NuWro. However, for one discussed case, we consider the newest version of NuWro, ver. 25.03. As we changed the boundary conditions, we repeated the previous analysis and obtained a new model for neutrino-carbon scattering. This model serves as the baseline for subsequent studies.
Figure 1: Generator architecture. The blue block encloses the portion of the network that remains frozen when training with TL.
Figure 2: Discriminator architecture. The blue block encloses the portion of the network that remains frozen when training with TL.
II.2 Networks
We consider the same network architecture as in our previous paper [23] (model for inclusive CC neutrino-carbon scattering). The generator takes as input a random vector of length $50$ (drawn from a Gaussian multivariate distribution) and neutrino energy $E^{\prime}_{\nu}$. Both inputs are processed independently by two distinct dense layers. Then, the output is concatenated and processed by three blocks, as shown in Fig. 1.
To obtain the GAN model, one needs to train the discriminator model, which takes as input the proxy variables $\theta^{\prime}$ and $E_{\mu}$, processed by two independent blocks. The output is then concatenated and transformed. Eventually, the output is transformed by four blocks, as shown in Fig. 2.
The fundamental building block of our network consists of a fully connected (dense) layer, followed by layer normalization [27], an activation function—ReLU[28] for the generator and PReLU[29] for the discriminator, and a dropout layer. We refer to this configuration as Block1. Introducing a skip connection from the input to the output of Block1 yields Block2. When the $E^{\prime}_{\nu}$ information is concatenated to the input of Block1 and subsequently Block2, we obtain Block3 and Block4, respectively. Furthermore, Gaussian noise is injected into both the generator and the discriminator after each layer containing learnable parameters. A comprehensive description of the GAN architecture is provided in Ref.[23].
Although the architectural scheme remains unchanged, we introduce several modifications compared to the previous analysis. Specifically, Gaussian noise is now injected into both the generator and the discriminator, whereas previously it was applied only for the discriminator. In addition, Blocks 1 and 2 in the main body of the generator have been replaced by Blocks 3 and 4, respectively, as illustrated in Fig. 1. It is important to note that the discriminator architecture remains unchanged (see Fig. 2). Notably, the discriminator is approximately twice as deep as the generator, and the number of neurons per layer is configured to yield roughly twice the number of trainable parameters per layer compared to the generator.
II.3 Loss and metrics
The loss function has two components: one for the discriminator and one for the generator. As in our previous study, the discriminator loss is binary cross-entropy. In the case of the generator loss, we tested three cases, the same choice as in our previous work, the “classical” one initially proposed in Ref. [15], the heuristic (non-saturating) loss also proposed in Ref. [15], and the sum of both. These preliminary studies showed that in the case of our analyses, the last proposal works the most efficiently. As it is not a common choice of the loss, we discuss it in the Appendix A.
The network models are optimized using the AdamW [30] algorithm in the initial stage of the training and then with Amsgrad [31].
Figure 3: Testdata samples $(E^{\prime}_{\mu},\theta^{\prime})$ distributions for $E_{\nu}=500$ MeV. On the top left, we have the original $\nu_{\mu}$-carbon, on the top right is the $\nu_{\mu}$-argon, on the bottom left, $\bar{\nu}_{\mu}$-carbon, and on the bottom right, the alternative $\nu_{\mu}$-carbon samples’ distributions.
Figure 4: Training data $450<E_{\nu}<550$ MeV sub-samples $(E^{\prime}_{\mu},\theta^{\prime})$ distributions. On the left side, we have the sub-sample from a training sample of 100100k events, while on the right side is the sub-sample from a sample of 1010k events.
To evaluate the quality of the model, we consider two metrics: the Mean Averaged Pull (MAP) [32, 18, 23] and the Earth Mover’s Distance (EMD) [33]. In our earlier studies, we used EMD in its two-dimensional form; however, in this analysis, we adapt the MAP to three dimensions (MAP-3D). This adaptation involves creating a histogram of the variables ($E^{\prime}_{\nu}$, $E^{\prime}_{\mu}$, $\theta^{\prime}$) using a uniformly distributed sample of $E_{\nu}$ with a binning of $14\times 14\times 14$.
For the MAP-3D metric, we consider two versions: one that includes all tails, where the bin content of both the NuWro histogram ($n_{nuw,i}\neq 0$) and the GAN histogram ($n_{gan,i}\neq 0$) is used, and another that excludes low-content bins, where we only consider bins with $n_{nuw,i}>5$ and $n_{gan,i}>5$. Additionally, we compute the MAP metrics using histograms of the two-dimensional distribution of $E_{\mu}^{\prime}$ and $\theta^{\prime}$, with a binning of $50\times 50$. The metric values during optimization are calculated on a validation dataset consisting of one million events.
Typically, the best model is the one that minimizes both metrics, but, in the present analysis, to choose the most optimal model, we evaluate an average of both MAP-3D metrics $\langle\text{MAP}\rangle=\frac{1}{2}(\text{MAP-3D}+\text{MAP-3D w/o tails})$. The selected model minimizes this metric.
Figure 5: Dependence of the MAP-3D metric on the number of optimization epochs for the $\nu_{\mu}$–argon models. The left and right columns correspond to optimizations using 10,000 and 100,000 events, respectively. The top and bottom rows show results for models trained from scratch and with transfer learning (TL), respectively. Red dashed vertical lines mark the epochs of the selected models.
III Transfer learning
The main idea of our approach is to use a pre-trained model (baseline model) that was trained on inclusive CC muon neutrino scattering off the carbon target. The new training data are uniformly distributed in the energy range ($300$ MeV, $10$ GeV), similar to our previous analysis.
Using the pre-trained model, we apply the concept of transfer learning [34, 35, 36], assuming that the generator and discriminator have already learned the fundamental patterns underlying neutrino–nucleus interactions. We fine-tune the pre-trained model on smaller training datasets than those used during its initial optimization. This approach leverages the prior knowledge encoded by the model, allowing it to adapt efficiently to new data. A similar strategy was successfully employed in our previous work on modeling inclusive electron scattering data [26].
One of the key advantages of the transfer learning technique is that it reduces the need for large amounts of new data when developing models for related processes. Since the pre-trained model already encodes fundamental relationships underlying neutrino scattering, only a relatively small amount of additional data is required to fine-tune the model to a new but related scattering scenario.
To adapt the pre-trained model to new data, we freeze the first block of both the generator and discriminator, as illustrated in Figs. 1 and 2. The remaining blocks are fine-tuned during the optimization process.
To verify the efficiency of the TL concept, we consider two training data sizes for each process: 100,000 and 10,000 events. Additionally, in each case, we train the model from scratch.
Table 1: Some of the performance metrics of the pre-trained $\nu_{\mu}$-carbon model.
$E_{\nu}$ EMD MAP (3D) MAP w/o tails (3D)
500 MeV 0.15 1.09 1.09
1 GeV 0.15 1.05 1.05
3 GeV 0.15 1.00 1.00
5 GeV 0.15 1.00 0.98
9 GeV 0.14 0.97 0.95
All 0.12 0.88 0.89 0.86 0.88
Table 2: Performance metrics for generators of CC inclusive $\nu_{\mu}$–argon scattering events. The left and right panels display results for the models trained from scratch and using transfer learning, respectively. The training dataset comprises 10,000 (top table) and 1000,000 (bottom table) events.
Trained on 10,000 events
Scratch TL
$E_{\nu}$ EMD MAP (3D) MAP w/o tails (3D) EMD MAP (3D) MAP w/o tails (3D)
500 MeV 0.96 6.46 7.63 0.36 2.51 2.71
1 GeV 0.84 5.42 6.17 0.37 2.42 2.67
3 GeV 0.73 3.92 4.30 0.25 1.40 1.44
5 GeV 0.65 3.50 3.81 0.23 1.32 1.35
9 GeV 0.65 3.58 3.88 0.28 1.60 1.65
All 0.36 2.01 3.01 2.05 3.22 0.17 1.07 1.27 1.08 1.29
Trained on 100,000 events
500 MeV 0.32 2.42 2.67 0.21 1.63 1.69
1 GeV 0.26 1.58 1.64 0.19 1.38 1.41
3 GeV 0.20 1.30 1.33 0.16 1.10 1.10
5 GeV 0.18 1.21 1.21 0.18 1.14 1.15
9 GeV 0.20 1.31 1.33 0.20 1.41 1.45
All 0.14 0.98 1.06 0.97 1.05 0.14 0.94 1.04 0.93 1.03
III.1 Domain adaptation
We consider three types of inclusive CC scattering processes:
  1. (A)
    $\nu_{\mu}$-argon scattering;
  2. (B)
    $\bar{\nu}_{\mu}$-carbon scattering;
  3. (C)
    $\nu_{\mu}$-carbon scattering but with a different model of interaction than used in the pre-training.
The pre-trained model generates events that follow the distributions predicted by the NuWro ver.21.09 simulation, which employs the spectral function approach to model the carbon (12C) nucleus in the quasielastic (QE) channel [37], and the local Fermi gas model for other interaction channels. The generated events may include a charged muon and one or more nucleons, as well as occasionally pions, resulting from final-state interaction (FSI) effects. Pion production is modeled solely through the $\Delta(1232)$ resonance [38, 39, 40], while contributions from more inelastic processes are modeled using parton distribution functions (PDFs).
In scenario (A), we consider the argon (40Ar) nucleus, described by a spectral function [41] derived from the analysis of Jefferson Lab coincidence electron–argon scattering data [42, 43]. This setup allows us to test the effectiveness of TL when the nuclear target is changed from carbon to a structurally different nucleus.
In Scenario (B), we use the same model for nuclear and hadronic interactions; however, instead of neutrino–carbon interactions, we consider muon antineutrino–carbon interactions.
Scenario (C) addresses a case where the default model used during pre-training encodes interaction dynamics that differ from those in the fine-tuning data. Specifically, the fine-tuning dataset is generated using the latest version of NuWro, ver.25.03, which includes an improved treatment of resonant pion production extended to invariant masses up to $1.9$ GeV. In addition, we deliberately modify key model parameters: the axial mass in the QE channel is increased from $1.030$ GeV to $1.350$ GeV, and the axial mass in the pion production model is changed from $0.94$ GeV to $1.034$ GeV. Finally, the nuclear model is altered by replacing the spectral function with the local Fermi gas model.
The three scenarios differ in their kinematic domains, displaying notable variations in coverage of the relevant phase space, as shown in Fig. 3. This figure presents the $(E^{\prime}_{\mu},\theta^{\prime})$ distributions for 1,000,000 test events in each case: the original test dataset ($\nu_{\mu}$–carbon), and the three test scenarios: $\nu_{\mu}$–argon, $\bar{\nu}_{\mu}$–carbon, and $\nu_{\mu}$–carbon with a modified interaction model. The four datasets exhibit significant differences in their distributions. The most pronounced divergence is observed between the (anti)neutrino–carbon cases and the two involving $\nu_{\mu}$–argon and the altered $\nu_{\mu}$–carbon interaction model. In the latter, a distinct band associated with the QE peak is visible.
Figure 6: Reconstructed hadronic invariant mass $W$-distributions, for inclusive CC $\nu_{\mu}$-argon scattering, generated by NuWro and GAN model trained from scratch and using TL, for the optimization with 10k events. We present the results for $E_{\nu}=500$ MeV (top left), 11 GeV (top right), 33 GeV (bottom left), and 55 GeV (bottom right).
Figure 7: Same as in Fig. 6 but for $C$ $\nu_{\mu}$-argon scattering and the training dataset with 100,000 events.
Table 3: Same as in Table 2 but for charged-current inclusive $\bar{\nu}_{\mu}$–carbon scattering.
Trained on 10,000 of events
Scratch TL
$E_{\nu}$ EMD MAP (3D) MAP w/o tails (3D) EMD MAP (3D) MAP w/o tails (3D)
500 MeV 0.64 4.65 5.21 0.36 2.13 2.32
1 GeV 0.88 4.45 5.00 0.38 2.57 2.79
3 GeV 0.84 4.00 4.46 0.39 2.07 2.21
5 GeV 0.67 3.57 3.87 0.33 1.67 1.75
9 GeV 0.72 3.77 3.98 0.32 1.85 1.94
All 0.51 2.15 2.88 2.37 3.10 0.21 1.26 1.56 1.28 1.61
Trained on 100,000 of events
500 MeV 0.31 2.27 2.41 0.22 1.63 1.66
1 GeV 0.29 1.77 1.79 0.25 1.52 1.55
3 GeV 0.20 1.39 1.40 0.18 1.24 1.25
5 GeV 0.21 1.24 1.24 0.20 1.20 1.21
9 GeV 0.25 1.33 1.34 0.22 1.24 1.26
All 0.13 0.93 1.10 0.91 1.09 0.14 0.94 1.08 0.92 1.08
pRefer to caption Refer to caption Refer to caption Refer to caption
Figure 8: Same as in Fig. 6 but for $C$ $\bar{\nu}_{\mu}$-carbon scattering and the training dataset with 10,000 events.
Figure 9: Same as in Fig. 6 but for $C$ $\bar{\nu}_{\mu}$-carbon scattering and the training dataset with 100,000 events.
Table 4: Performance metrics for generators of $C$ $\nu_{\mu}$–carbon (alternative interaction model) scattering events. The left and right panels display results for the models trained from scratch and using transfer learning, respectively. The training dataset comprises 10,000.
Scratch TL
$E_{\nu}$ EMD MAP (3D) MAP w/o tails (3D) EMD MAP (3D) MAP w/o tails (3D)
500 MeV 1.18 8.47 10.20 0.51 3.48 3.99
1 GeV 1.05 6.27 7.32 0.44 2.95 3.25
3 GeV 1.03 5.14 5.83 0.35 2.16 2.28
5 GeV 0.95 4.40 4.86 0.35 2.13 2.23
9 GeV 0.86 4.46 4.86 0.35 2.12 2.22
All 0.47 2.58 3.54 2.64 3.86 0.20 1.23 1.63 1.23 1.68
Figure 10: Same as in Fig. 6 but for $C$ $\nu_{\mu}$-carbon (alternative interaction) scattering and the training dataset with 10,000 events.
IV Numerical results
To study the effectiveness of transfer learning, we initiated our work by obtaining a baseline (pre-trained) generator and discriminator for inclusive neutrino-carbon scattering events. We trained the model on samples uniformly distributed across different neutrino energies (from $300$ MeV to $10$ Gev), using a training dataset comprising four million events. To test the quality of this model, we generated new data at fixed neutrino energies: $500$ MeV, $1$ GeV, $3$ GeV, $5$ GeV, and $10$ GeV, as well as for uniformly distributed samples across the entire energy spectrum. The values of performance metrics of the baseline GAN model are low and comparable to those reported in our previous paper. Table 1 summarizes the quality of the baseline model.
To evaluate the effectiveness of the transfer learning technique – specifically, to determine how well the baseline model captures universal patterns shared across all processes discussed in this paper – we consider two training scenarios. In particular, we train models using both a large dataset of 100,000 and a smaller one containing 10,000 events.
Fig. 4 illustrates the differences in event distributions between the two training datasets for $\nu_{\mu}$–argon scattering events within the neutrino energy range of $450$ to $550$ MeV. The figure clearly shows that training the GAN and discriminator with only 10,000 events poses a significant challenge.
We begin our discussion with case (A): neutrino scattering on an argon target.
To highlight the differences in optimization between training on 10,000 and 100,000 events, Fig. 5 shows the evolution of performance metrics during training. For the training with 100,000 events, the model’s performance remains stable. It is noteworthy that the best-performing models in the transfer learning framework are reached significantly faster than those trained from scratch. Furthermore, as summarized in Table 2, the transfer learning models consistently achieve lower metric values, indicating better performance. This suggests that when trained from scratch, the network struggles to learn the features of $\nu_{\mu}$–argon scattering from limited data, whereas the transfer learning approach, which builds on prior knowledge of $\nu_{\mu}$–carbon interactions, offers a substantial advantage. Eventually, when the number of samples in the training dataset increases, the quality of the obtained GAN models improves. However, even with a large number of samples, the transfer learning model still outperforms the model trained from scratch.
Figs. 6 and 7 show the event distributions in reconstructed hadronic invariant mass $W$ generated by our GAN models and NuWro for training datasets containing 10,000 and 100,000 events, respectively. This quantity is defined by $W=\sqrt{M_{N}^{2}+2\omega M-Q^{2}}$, with $\omega=E_{\nu}-E_{\mu}$ the energy transferred by the neutrino to the struck nucleus, $M$ the average nucleon mass, and $Q^{2}$ the four-momentum transfer squared $Q^{2}=2E_{\nu}(E_{\mu}-p_{\mu,z})-m_{\mu}^{2}$. The $W$ distributions clearly highlight the quasi-elastic and $\Delta(1232)$ resonance peaks. We show simulations for several neutrino energies: $500$ MeV, $1$, 33, and $5$ GeV. As illustrated, the GAN model trained from scratch struggles to capture the correct structure of the distributions across all energy values. In particular, Fig. 6 highlights the model’s failure to reproduce the peak in the $W$ distribution, whereas the model pre-trained on carbon data achieves a remarkably accurate fit. While using a larger, more informative training dataset improves the performance of the model trained from scratch, it still does not reach the accuracy achieved through transfer learning.
Similarly to case (A), we perform an analogous analysis for case (B), involving $\bar{\nu}_{\mu}$–carbon interactions. The results are summarized in Table 3. Models trained using transfer learning consistently outperform those trained from scratch. When trained on a larger and more informative dataset, model performance improves significantly.
Figs. 8 and 9 show histograms generated by our models and by NuWro for energies of $500$ MeV, and $1$, 33, and $5$ GeV. In contrast to the neutrino–argon case, even the models trained from scratch can capture the resonance structures present in the distributions. Nevertheless, models trained with transfer learning demonstrate superior efficiency and performance.
In the final case (C), we optimize two models, one trained from scratch and the other using transfer learning, using only a small training dataset. Table 4 and Fig. 10 summarize our results. The model trained from scratch struggles to accurately capture the QE peak and performs poorly in the resonance region. In contrast, the model trained via transfer learning performs significantly better, despite being initially trained on data generated by a completely different NuWro model.
V Summary
We have trained GAN models using transfer learning from a pre-trained GAN originally developed on the inclusive CC $\nu_{\mu}$–carbon MC dataset, adapting them to various other scattering scenarios, including inclusive CC $\bar{\nu}_{\mu}$–carbon, $\nu_{\mu}$–argon, and $\nu_{\mu}$–carbon with alternative interaction models. These models were compared against their counterparts trained from scratch. We observed that initializing from the pre-trained model consistently outperforms training from scratch, particularly when datasets have limited statistical significance or provide a poor sampling of the kinematic distributions of the underlying physical processes. This suggests that the pre-trained model captures essential features of neutrino scattering dynamics.
To evaluate the effectiveness of transfer learning more thoroughly, we also trained the same network architecture on datasets with a significantly larger number of events. This demonstrated that the inferior performance of models trained from scratch on smaller datasets is not due to architectural limitations. In this larger-statistics scenario, the model trained from scratch begins to catch up with the transfer-learned model. Nevertheless, the latter continues to show advantages, particularly in regions where the dataset remains a poor representation of the relevant kinematic phase space.
These findings highlight the promising potential of transfer learning in developing more accurate models for (anti)neutrino–nucleus scattering. In particular, one can start from a model trained on MC simulation data and adapt it to more realistic scenarios using comparatively limited data. This approach offers a practical path forward, especially in cases where experimental statistics are sparse. Further investigation in this direction is planned for future work.
Appendix A Loss function
Figure 11: Kinematic distributions for inclusive CC $\nu_{\mu}$-carbon scattering for $E^{\prime}_{\mu}$ and $\theta^{\prime}$ variables computed for 1M events at different neutrino energies, according to NuWro. From left to right and top to bottom distributions for energies $E_{\nu}=$ $0.35$, $0.5$, 11, 33, 55, and $9$ GeV.
Figure 12: Probability distributions from the discriminator for samples generated by NuWro and by the GAN trained with the heuristic loss $L_{H}$ [Eq. (4)] and with our loss $L_{HCE}$ [Eq. (6)], for $E_{\nu}=350$ MeV. Here, NN and $G$ denote samples from NuWro and the GAN, respectively.
Typically, in most applications, the heuristic non-saturating loss
$L_{H}(G)=-\frac{1}{B}\sum_{k=1}^{B}\log[D(G(\mathbf{x}k,E{\nu,k}),E_{\nu,k})]$ (4)
is considered, where $G$ and $D$ denote the generator and discriminator, respectively; $B$ is the mini-batch size; $\mathbf{x}_{k}$ is the latent space vector; and $E_{\nu,k}$ represents the energy of the interacting neutrino.
The rationale for using the loss (4) stems from the fact that the maximization of the conventional cross-entropy loss,
$L_{CE}(G)=\frac{1}{B}\sum_{k=1}^{B}\log[1-D(G(\mathbf{x}k,E{\nu,k}),E_{\nu,k})],$ (5)
tends to saturate early in the training process, as the discriminator quickly becomes adept at distinguishing generated events from real ones. Nonetheless, $L_{H}$ has a saturation region as well, just in the “opposite side” compared to $L_{CE}$.
However, we observe that employing the non-saturating loss in Eq.(4) can lead to uneven performance of the discriminator across different $E_{\nu}$ regions. Specifically, the GAN appears to favor optimization at higher neutrino energies, where the model benefits from smoother kinematic variations across adjacent energy bins. In contrast, at lower energies, slight variations in $E_{\nu}$ produce significant changes in the corresponding kinematic distributions, making training more challenging. The kinematic domains covered by the data for various neutrino energies are illustrated in Fig. 11. At low $E_{\nu}$ energies, $350$ and $500$ MeV, mostly the quasi-elastic contribution is visible. In contrast, at higher energies, resonance contributions also play a significant role. Since higher-energy events dominate our training data, the models being trained first capture the structures associated with these higher energies.
Figure 13: Same as Fig. 12, but for $E_{\nu}=5$ GeV.
Figure 14: Kinematic distributions, pull per bin, and pull distributions for samples from NuWro and GANs trained with the heuristic loss, for $E_{\nu}=500$ MeV (first two rows) and $E_{\nu}=3$ GeV (last two rows).
Figure 15: Same as in Fig. 14 but for model optimized with our loss (5).
A special role in the optimization plays the discriminator, whose ability to distinguish between “fake” and “true” events is essential for obtaining a good generator. If the discriminator works either perfectly or imperfectly, then one can not get a good GAN model.
For an optimal model of GAN, one can expect the response of the discriminator, which is a probability given by the sigmoid function $Sigmoid(D)={1}/{1+\exp(D)}$, to be distributed around the value $0.5$. Such distributions are shown in Figs. 12 and 13, for low and higher neutrino energy events with $E_{\nu}=350$ MeV and $E_{\nu}=5$ GeV, respectively. These figures display the discriminator distributions for NuWro (“true”) and generated by the GAN distribution of events. The GAN was trained with heuristic loss $L_{H}$ defined in Eq. 4. We observe that the models trained with the heuristic loss exhibit greater dispersion, in low neutrino energy, in their predictions, whereas at higher energies, their distributions are more tightly constrained. Moreover, the lower energy distribution is peaked at a value lower than $0.5$ in contrast to the higher energy distribution. These deficiencies of the discriminator, in low neutrino energies, lead to inefficiency in the optimization of the GAN model.
It turns out that to optimize efficiently for both low and high neutrino energy cases, it is sufficient to consider the following loss function:
$L_{HCE}(G)$ $=$ $\frac{1}{B}\sum_{k=1}^{B}\left\{\log\left[1-D(G(\mathbf{x}_{k},E_{\nu,k}),E_{\nu,k})\right]\right$. (6)
$\left.-\log\left[D(G(\mathbf{x}_{k},E_{\nu,k}),E_{\nu,k})\right]\right\},$
which combines the original non-saturating heuristic loss (4) with the maximization term from the conventional cross-entropy (5).
The distributions of discriminator probabilities, obtained with (6), are also shown in Figs. 12 and 13. The pathological shape of the discriminator distribution at low energies disappears, and it is peaked at the value $0.5$.
Additionally, in Fig. 14 and 15 we show 2-dimensional distributions of proxy variables for low and higher neutrino energies ($E_{\nu}=0.5$ and $E_{\nu}=3$ GeV) generated by NuWro and GAN models obtained with heuristic and proposed loss by us, respectively. We observe that at lower neutrino energies, the GAN model trained with heuristic loss fails to reproduce a fraction of events on the boundary of the allowed region.
Figure 16: Dependence on $W$ of the discriminator-predicted probabilities for events generated by GANs trained with the loss functions in Eqs. (4) (left) and (5) (right), for $E_{\nu}=350$ MeV.
Figure 17: Same as in Fig. 16 but for $E_{\nu}=5$ GeV.
Eventually, in Figs. 16 and 17, we show a comparison between $W$-dependence of probability distributions predicted by discriminators for the events generated by the GANs obtained using (4) and (5) loss functions. We observe that the model optimized with the loss proposed in this paper exhibits a distribution concentrated around the probability $0.5$ for both presented neutrino energies.
In contrast, for the model with heuristic loss for low neutrino energy, Fig. 16, besides the events concentrated around probability $0.5$, there are two groups distributed around probability values $0.45$ and $0.55$ for $W$ about $0.85$ and $1.05$ GeV, respectively. These two regions correspond to quasi-elastic and $\Delta$ resonance production contributions.
For higher neutrino energies, Fig. 17, the $W$-distributions for both models are very similar.
Acknowledgements.
This research is partly (K.M.G., A.M.A., J.T.S.) or fully (B.E.K., J.L.B, H.P., R.D.B.) supported by the National Science Centre under grant UMO-2021/41/B/ST2/ 02778. J.L.B. and J.T.S. are also supported by the Polish Ministry of Science Grant No. 2022/WK/15. K.M.G is partly supported by the “Excellence Initiative – Research University” for the years 2020-2026 at the University of Wrocław.